Using the Chain rule find the derivative of the function: 1. y=(x?+2x2,6
Answer #1

Given function ,  y=(x^{3}+2x^{2})^{6}

Differentiating both side with respect to  x we get ,

\frac{dy}{dx}=\frac{d}{dx}(x^{3}+2x^{2})^{6}

\Rightarrow \frac{dy}{dx}=6(x^{3}+2x^{2})^{5}\frac{d}{dx}(x^{3}+2x^{2}) , Using Chain rule .

\Rightarrow \frac{dy}{dx}=6(x^{3}+2x^{2})^{5}\left \{ \frac{d}{dx}(x^{3})+2\frac{d}{dx}(x^{2}) \right \}

\Rightarrow \frac{dy}{dx}=6(x^{3}+2x^{2})^{5}(3x^{2}+4x)

Hence ,

Answer :  \frac{dy}{dx}=6(x^{3}+2x^{2})^{5}(3x^{2}+4x)